OpenStudy (anonymous):

PLZZZZ HELP!!!! 3. You recently started the paperwork to purchase your new home, ad were just notified that you can move into the house in 2 weeks. You decide to hire a moving company, but are unure which company to choose. You seach online and are interested in contacting two companies, Heavy Liting and Quick Move, to discuss their rates. Heavy Lifting charges a $80 fee plus $35 per hour. Quick Move charges $55 per hour with no additional fees.

OpenStudy (anonymous):

a. which mover provides a better deal for 2 hours of work? How did you arrive at your answer? quick move is cheaper @ $110 compared @ $150 b.Which mover provides a better deal for 15 hours of work? How did you arrive at your answer? 825 better deal is Heavy Lifting c. For what values h (hours) doseQuick Move offer the better deal? Express your answer as an inequality. Explain how you reached you answer.

OpenStudy (anonymous):

you have the answer for a) yes?

OpenStudy (anonymous):

also b)

OpenStudy (anonymous):

hold on that was wrong. quick move is better if the hours are low, heavy lifting if the hours are big

OpenStudy (anonymous):

put \[55x=80+35x\] and solve for x

OpenStudy (anonymous):

\[55x=80+35x\] \[20x=80\] \[x=4\] so they are the same for 4 hours of work

OpenStudy (anonymous):

more than 4 hours heavy lifting is cheaper. less that four hours quick move is

OpenStudy (anonymous):

THANK YOU THANK YOU THANK YOU!!!!!